OpenStudy (anonymous):

Find an equation of the tangent line to the graph of f at the given point. f(x) = x^2 + 3, (1, 4)

OpenStudy (tkhunny):

Have you considered the 1st derivative, f'(x) evaluated at x = 1 to give the slope of this mysterious tangent line?

OpenStudy (anonymous):

I don't know how to do this problem.

OpenStudy (tkhunny):

Have you heard of the 1st Derivative?

OpenStudy (anonymous):

no

OpenStudy (tkhunny):

Is this a placement test?

OpenStudy (anonymous):

no, homework. My teacher is foreign & hard to understand

OpenStudy (tkhunny):

What course is this? Is it Calculus I?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

if you can give me an equation, I can probably figure it out, just don't know what the (

OpenStudy (tkhunny):

Bear with me. Just trying to figure out where you are. Have you been introduced to an odd-looking ratio, like this \(\dfrac{f(x+h) - f(x)}{h}\)?

OpenStudy (anonymous):

sorry I lost you. I know that equation, just don't know what the (1,4) is for?

OpenStudy (tkhunny):

The very first thing to do is to verify that (1,4) is actually ON the curve. Is it?

OpenStudy (anonymous):

I don't think so

OpenStudy (anonymous):

not totally sure how to read the graphs either

OpenStudy (tkhunny):

f(x) = x^2 + 3 f(1) = 1^2 + 3 = 1+3 = 4 (1,4) is on f(x) You ARE going to have to get better at some of these basic skills. This class is going to be a LONG HARD struggle if you don't remember anything from your prerequisites. Algebra I Algebra II Geometry Trigonometry Hopefully Analytic Geometry Hopefully "College Algebra" We're going to need all the algebra and geometry skills we can get to survive just today. What do you think? Are you up to it?

OpenStudy (anonymous):

I am. This is my last math class, so I do need to pass it. Thank you!

OpenStudy (tkhunny):

Okay, then... In the Differential Calculus, we have this amazing thing known as the 1st Derivative. It can be loosely defined as the slope of the tangent line at any given point on a curve (where it works, anyway). Our task is to define a little more carefully exactly what that means. You may be familiar with a line tangent to a circle. It is a line that just touches the circle. There is only one point of intersection. If you hit the circle twice, it is a secant line. How are we doing? Up to speed on these concepts?